# DiaSynth: Synthetic Dialogue Generation Framework for Low Resource Dialogue Applications

### v2

@@ Abstract
The scarcity of domain-specific dialogue datasets limits the development of dialogue systems across applications. Existing research is constrained by general or niche datasets that lack sufficient scale for training dialogue systems. To address this gap, we introduce DiaSynth - a synthetic dialogue generation framework capable of generating high-quality, contextually rich dialogues across a wide range of domains. Unlike existing frameworks, DiaSynth uses Large Language Models (LLMs) and Chain of Thought (CoT) reasoning to generate dynamic, domain-specific dialogues with simulated personas and diverse conversational features. We perform our experiments by generating synthetic data using different LLMs and few-shot examples from DialogSum and SAMSum. The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47% on dialogue summarization, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the performance distribution of the in-domain data on dialogue summarization. The quality of the data generated also increases as we increase the size of LLM from 3B to 8B. These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods. We open source the code and data generated for future research.
@@ 3.2 Persona Generation
Personas of the individuals involved in a conversation are primary influencers in determining how a conversation pans out. Using random personas from persona datasets and prompting the LLM to simulate a dialogue between them about a random topic often leads to superficial dialogues that lack depth and contextual richness. To address this issue, DiaSynth generates a user-specified number of personas for each subtopic, ensuring that the personas are conditioned on the subtopic context. This conditioning prompts the LLM to create personas that are most likely to engage in a meaningful dialogue about the subtopic, such as a medical professional and a patient discussing "medical diagnostics" or a researcher and a student talking about "academic publishing." We also run a similarity check for the personas too. The conditioned persona generation is crucial because it ensures that future dialogues will not only be contextually rich but also exhibit a high level of depth. Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting.
@@ 4.2 Downstream Task - Summarization
We fine-tune pretrained summarization models like DistilBART, BART Lewis et al. (2020), T5 Raffel et al. (2020) and LED Beltagy et al. (2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics. We evaluate the usability of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
@@ 5 Results
This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs and varying few-shot examples. Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets. These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples. In total, eight distinct datasets were generated using DiaSynth by pairing each LLM with the two sets of few-shot examples, resulting in all possible combinations. For each combination, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic. The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 1. All the experiments were run on a single A100 GPU with the generation time ranging from 2 hours to 4 hours.
@@ Appendix A Hallucination Study
In addition to evaluating the quality and usability of dialogues produced by DiaSynth, we conducted a study on the phenomenon of hallucinations within the generated dialogues. Hallucinations in language models refer to instances where the output contains misleading or incorrect information or situations where the model repeats the same content. To evaluate the occurrence of hallucinations, we compared the generated dialogues with their respective summaries and assessed them using two well-known hallucination benchmarks: SelfCheckGPT Manakul et al. [2023] and ChainPoll Friel and Sanyal [2023]. This analysis provides insights into the prevalence of hallucinations and informs strategies for improving dialogue quality in future iterations of DiaSynth. The results are presented in Tables 10 and 10.
@@ Appendix B Scalability of DiaSynth
1def calculate_total_dialogs(n, m, p):
2 """
3 Calculate the total number of dialogs generated.
4
5 Parameters:
6 n (int): Number of topics
7 m (int): Number of subtopics per topic
8 p (int): Number of personas per subtopic
9
10 Returns:
11 int: Total number of dialogs generated
12 """
13 # Calculate the total number of subtopics
14 total_subtopics = n * m
15
16 # Calculate the number of dialogs for each subtopic using combinations of personas
17 dialogs_per_subtopic = (p * (p - 1)) // 2
18
19 # Total dialogs generated across all subtopics
20 total_dialogs = total_subtopics * dialogs_per_subtopic
21
22 return total_dialogs
Listing 1: Calculation of Total Dialogs Generated
@@ Appendix D Example CoT environments
(a) CoT Example 1
(b) CoT Example 2
Figure 2: Illustration of Example CoT environments
@@ Appendix E Ablation study on the use of CoT
In this section of the appendix, we present the impact of CoT in the dialogues generated by DiaSynth. We generate two datasets without CoT, using Phi-3 using DialogSum and SAMSum as few-shot examples with 8 topics. Tables 13, 13, 15 and 15 compare the scores of dialogues generated with and without CoT for the FED score and GPTScore and it can be clearly that CoT tends to increase the quality of the dialogues generated by DiaSynth.

### v3

@@ Abstract
The scarcity of domain-specific dialogue datasets limits the development of dialogue systems across applications. Existing research is constrained by general or niche datasets that lack sufficient scale for training dialogue systems. To address this gap, we introduce DiaSynth - a synthetic dialogue generation framework capable of generating high-quality, contextually rich dialogues across a wide range of domains. Unlike existing frameworks, DiaSynth uses Large Language Models (LLMs) and Chain of Thought (CoT) reasoning to generate dynamic, domain-specific dialogues with simulated personas and diverse conversational features. We perform our experiments by generating synthetic data using different LLMs and few-shot examples from DialogSum and SAMSum. The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47% on dialogue summarization, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the performance distribution of the in-domain data on dialogue summarization. The quality of the data generated also increases as we increase the size of LLM from 3B to 8B. These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods. [^1] https://github.com/ntuspeechlab/DiaSynth We open source the code and data generated for future research.
DiaSynth: Synthetic Dialogue Generation Framework for Low Resource Dialogue Applications
Sathya Krishnan Suresh1, Wu Mengjun1, Tushar Pranav2, Eng Siong Chng1 1Nanyang Technological University 2Singapore Institute of Technology Correspondence: sathyakr001@e.ntu.edu.sg
@@ 3.2 Persona Generation
Personas of the individuals involved in a conversation are primary influencers in determining how a conversation pans out. Using random personas from persona datasets and prompting the LLM to simulate a dialogue between them about a random topic often leads to superficial dialogues that lack depth and contextual richness. To address this issue, DiaSynth generates a user-specified number of personas for each subtopic, ensuring that the personas are conditioned on the subtopic context. This conditioning prompts the LLM to create personas that are most likely to engage in a meaningful dialogue about the subtopic, such as a medical professional and a patient discussing "medical diagnostics" or a researcher and a student talking about "academic publishing." We also run a similarity check for the personas too. The conditioned persona generation is crucial because it ensures that future dialogues will not only be contextually rich but also exhibit a high level of depth. Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting. We present the impacts of sub-topics and personas in Appendix F
@@ 4.2 Downstream Task - Summarization
We fine-tune pretrained summarization models like DistilBART, BART Lewis et al. (2020), T5 Raffel et al. (2020) and LED Beltagy et al. (2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics. We evaluate the usability of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data. We also present the results of finetuning BART and T5 on synthetic data and in-domain data on response generation in Appendix H.
@@ 5 Results
This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs and varying few-shot examples. Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets. These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples. In total, eight distinct datasets were generated using DiaSynth by pairing each LLM with the two sets of few-shot examples, resulting in all possible combinations. For each combination, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic. The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 1. All the experiments were run on a single A100 GPU with the generation time ranging from 2 hours to 4 hours. Additional methodolgical details are presented in Appendix G.
@@ Appendix A Hallucination Study
In addition to evaluating the quality and usability of dialogues produced by DiaSynth, we conducted a study on the phenomenon of hallucinations within the generated dialogues. Hallucinations in language models refer to instances where the output contains misleading or incorrect information or situations where the model repeats the same content. To evaluate the occurrence of hallucinations, we compared the generated dialogues with their respective summaries and assessed them using two well-known hallucination benchmarks: SelfCheckGPT Manakul et al. (2023) and ChainPoll Friel and Sanyal (2023). This analysis provides insights into the prevalence of hallucinations and informs strategies for improving dialogue quality in future iterations of DiaSynth. The results are presented in Tables 9 and 10.
@@ Appendix B Scalability of DiaSynth
Algorithm 1 Calculation of Total Dialogs Generated
1:procedure CalculateTotalDialogs(n, m, p)
[…]
5:     return total_dialogs
6:end procedure
@@ Appendix E Ablation study on the use of CoT
In this section of the appendix, we present the impact of CoT in the dialogues generated by DiaSynth. We generate two datasets without CoT, using Phi-3 using DialogSum and SAMSum as few-shot examples with 8 topics. Tables 12, 13, 14 and 15 compare the scores of dialogues generated with and without CoT for the FED score and GPTScore and it can be clearly that CoT tends to increase the quality of the dialogues generated by DiaSynth.
@@
Appendix F Ablation Studies on Subtopics and Personas
To further validate the effectiveness of the DiaSynth framework, we conducted ablation studies by evaluating the impact of removing sub-topics and personas from the data generation pipeline. The goal was to assess their contribution to the quality of the generated dialogues.
For these experiments, we generated approximately 960-1000 dialogues using Phi-3 and compared three settings:
  • subtopics: Removing sub-topics while keeping personas.
  • personas: Removing personas while keeping sub-topics.
  • diasynth: The full DiaSynth-generated data with both personas and sub-topics.
Metric subtopics personas diasynth
Coherent 0.9252 0.9584 0.9536
Error Recovery 0.9022 0.9414 0.944
Consistent 0.9095 0.9512 0.954
Diverse 0.9139 0.9512 0.9534
Depth 0.9193 0.9533 0.9521
Likeable 0.0069 0.007 -0.0005
Understandable 0.8918 0.9339 0.9353
Flexible -0.0038 -0.0042 0
Informative 0.0096 0.0072 0.009
Inquisitive 0.0129 0.0063 -0.0033
Table 16: FED scores with DialogSum as base
Metric subtopics personas diasynth
Coherence 0.0118 0.0098 0.0286
Diversity 0.0246 0.0198 0.031
Flexibility 0.0239 0.0121 0.0193
Understandable 0.0137 0.0121 0.0363
Inquisitive 0.0358 0.0308 0.0363
Consistent 0.0175 0.0132 0.0369
Informative 0.0242 0.0155 0.0201
Likeability 0.0083 0.0076 0.0213
Depth 0.0216 0.0181 0.0117
Error Recovery 0.0337 0.0249 0.0342
Table 17: GPTScore with DialogSum as base
Metric subtopics personas diasynth
Coherence 0.0106 0.0109 0.0325
Diversity 0.0224 0.0191 0.0372
Flexibility 0.0235 0.0141 0.0395
Understandable 0.0159 0.0252 0.0415
Inquisitive 0.0334 0.0137 0.0415
Consistent 0.0192 0.0415 0.037
Informative 0.0228 0.0155 0.021
Likeability 0.0128 0.0075 0.0232
Depth 0.0187 0.0158 0.0126
Error Recovery 0.0323 0.0213 0.029
Table 18: FED scores with SAMSum as base
Metric subtopics personas diasynth
Coherence 0.0106 0.0109 0.0325
Diversity 0.0224 0.0191 0.0372
Flexibility 0.0235 0.0141 0.0395
Understandable 0.0159 0.0252 0.0415
Inquisitive 0.0334 0.0137 0.0415
Consistent 0.0192 0.0415 0.037
Informative 0.0228 0.0155 0.021
Likeability 0.0128 0.0075 0.0232
Depth 0.0187 0.0158 0.0126
Error Recovery 0.0323 0.0213 0.029
Table 19: GPTScore with SAMSum as base
The ablation studies presented in Tables 16, 17, 18 and 19 demonstrate that the inclusion of both personas and sub-topics significantly enhances the quality of generated dialogues across FED and GPTScore metrics. For both DialogSum and SAMSum few-shot examples, dialogues generated with the full DiaSynth framework, incorporating both personas and sub-topics, achieved the highest scores in coherence, diversity, and consistency. This indicates that structured dialogue generation benefits from incorporating diverse sub-topics while maintaining persona-driven consistency.
An interesting observation arises in Table 18, where for the SAMSum dataset, the best-performing configuration involved using only sub-topics without personas. This deviation can likely be attributed to the more informal nature of SAMSum dialogues, where structured personas introduce a formal communication style that does not align well with the dataset. In contrast, in more structured datasets like DialogSum, the addition of personas provides clear improvements, ensuring dialogue coherence and natural flow.
Moreover, the largest improvements in quality are seen in coherence, error recovery, and understandability, particularly when both sub-topics and personas are included. While sub-topics alone contribute significantly to improving diversity and depth, their combination with personas enhances overall dialogue quality. This suggests that dataset characteristics play a crucial role in determining the effectiveness of persona modeling, highlighting the need for adaptive strategies in synthetic dialogue generation. Ultimately, these findings reinforce that DiaSynth-generated dialogues are robust and adaptable, providing high-quality synthetic data across both structured and informal conversational settings.
Appendix G Additional Methodological Details
G.1 Prompts for Dialogue and Summarization Generation
The prompts used for dialogue generation were carefully designed to guide the model in producing contextually rich and persona-driven conversations. The structure ensures that dialogues exhibit natural human-like characteristics, considering aspects such as speaker familiarity, emotional states, and formality levels.
For dialogue generation, the following detailed prompt was used:
You are an expert dialog generator. The following are examples of real-life dialogues.
example 1
"$dialogue"-"{dialogue_{1}}"$
example 2
"$dialogue"-"{dialogue_{2}}"$
Use the examples as references and generate a dialogue between people with the following personas:
$persona1-"{persona_{1}}"$
$persona2-"{persona_{2}}"$
Characteristics of the dialogue - to be understood before generating the dialogue: characteristics
- Assume the values for the characteristics and provide an explanation for choosing those values.
- These characteristics should be well understood as they implicitly affect and guide the conversation.
Chain of Thought Reasoning:
- Before generating the dialogue, reason about the values for each characteristic listed above.
- Explain in detail how each characteristic will influence a hypothetical dialogue between persons with those characteristics and personas.
- Ensure that the reasoning considers the interactions between different characteristics (e.g., how familiarity and emotional state might interact).
- This reasoning and explanation must be included between <cot> and </cot> tags. Do not skip this step.
- The dialogue generated should be based on the explanation provided.
This prompt ensures that the model generates structured and realistic dialogues by incorporating a reasoning step before dialogue generation. The inclusion of CoT reasoning forces the model to explicitly consider multiple conversational attributes, leading to more coherent and contextually appropriate responses.
This structured prompting approach enables the generation of high-quality dialogues. The complete list of prompts, including variations for different tasks, is available in our public code repository.
G.2 Seed Topic Selection
The seed topics were selected to ensure diversity across conversational scenarios. The topics used in our experiments are as follows:
Remote Work, Book Recommendations, Fitness Routines, Travel Destinations, Career Development, Movie Reviews, Video Games, Pets, Stock Market, Fashion Trends, Online Education, Mental Wellness, Climate Change, Sports, Artificial Intelligence, Food.
These topics were chosen from a mix of real-world discussion trends and common conversational themes, ensuring broad coverage and relevance for synthetic dialogue generation.
G.3 Hyperparameters for Generation
The hyperparameter settings for different stages of generation are reported below:
Subtopic Generation:
Temperature: 0.1 and max tokens: 2048
Persona Generation:
Temperature: 0.1 and max tokens: 2048
Dialogue Generation:
Temperature: 0.2 and max tokens: 4096
Appendix H Additional Downstream Task: Response Generation
Model Before Fine-tuning In-domain Data Llama3 GPT-4o
t5-base 0.4003 0.6870 0.6572 0.6612
bart-base 0.5681 0.6875 0.6721 0.6630
Table 20: BERTScore evaluation of response generation models fine-tuned on synthetic and in-domain datasets.
To further evaluate the utility of DiaSynth-generated data, we conducted an additional downstream task: response generation. This task was included to validate the effectiveness of our synthetic data beyond summarization. For these experiments, we selected Llama3 and GPT-4o as base datasets since they demonstrated superior performance across both quality and summarization metrics. We employed BERTScore as our evaluation metric due to its effectiveness in measuring the similarity between generated and reference responses. The results are presented in Table 20.
Both models show a significant improvement in BERTScore after fine-tuning on DiaSynth-generated data compared to the pre-trained baseline, highlighting its effectiveness in enhancing model performance. Additionally, models fine-tuned on DiaSynth data achieve scores that are close to those fine-tuned on in-domain data. For instance, t5-base achieves a BERTScore of 0.6572 on Llama3-generated data and 0.6612 on GPT-4o-generated data, compared to 0.6870 for in-domain fine-tuning. These results indicate that DiaSynth-generated data serves as a viable alternative for fine-tuning response generation models, performing comparably to in-domain data, even in low-resource scenarios.